# TLRN: Temporal Latent Residual Networks For Large Deformation Image Registration

### v2

@@ 4 Experimental Evaluation
We validate our model, TLRN, on both synthetic data and real CMR image videos. All our experiments are trained on servers with AMD EPYC $7502$ CPU of $126$GB memory and Nvidia GTX $4090$Ti GPUs. We use the Adam optimizer [15] with $20000$ epochs, learning rate of $1e^{-4}$, and a batch size of $32$.
@@ 4.1 Datasets
2D Lemniscate Sequence. We simulate a number of $1400$lemniscate" series ($64^{2}\times T$), where $T+1=12$ (examples are shown in Fig. 2). For each reference image $I_{i}^{0}$, we first generate a 2D contour with its coordinates computed by $\{^{x=a\cdot\text{cos}(\alpha)/(\text{sin}^{2}(\alpha)+1)}_{y=a\cdot\text{sin}(\alpha)\text{cos}(\alpha)/(\text{sin}^{2}(\alpha)+1)}$, where $a$ is a scaling factor, and $\alpha$ is uniformly sampled from $[0,2\pi]$. We then determine the contour thickness by varying the coordinates with parameters $\sigma_{x}$ and $\sigma_{y}$, i.e., $[x\pm\sigma_{x},y\pm\sigma_{y}]$. The subsequent time frames $\{I_{i}^{1},\cdots,I_{i}^{T}\}$ are deformed versions of the reference image, simulated by applying affine transformations such as scaling, rotation, and translation. We split the dataset into $1000/200/200$ sequences for training, validation, and testing.
@@ 4.2 Experimental Design
We assess the registration accuracy on CMR videos by performing registration-based segmentation of the left ventricular (LV) myocardium. The propagated segmentations by deforming manually segmented myocardium on the reference image using predicted transformation fields from all methods are then compared. To evaluate volume overlap between the propagated segmentation $A$ and the manual segmentation $B$, we compute the dice similarity coefficient (DSC) [7] by DSC$(A,B)=2(|A|\cap|B|)/(|A|+|B|)$, where $\cap$ denotes an intersection of two regions. Additionally, we measure the maximum discrepancy between boundaries of the propagated segmentation and the manual segmentation by Hausdorff distance (HD) [13]. Given two sets of boundary points $X\in A$ and $Y\in B$ for LV myocardium, we compute the distance by HD$(X,Y)=max\{h(X,Y),h(Y,X)\}$, where $h(X,Y)=\mathop{max}_{x_{i}\in X}\mathop{min}_{y_{i}\in Y}\|x_{i}-y_{i}\|$, and vice versa.

### v3

@@ 4 Experimental Evaluation
We validate our model, TLRN, on both synthetic data and real CMR image videos. All our experiments are trained on servers with AMD EPYC $7502$ CPU of 126126GB memory and Nvidia GTX 40904090Ti GPUs. We use the Adam optimizer [15] with $20000$ epochs, learning rate of $1e^{-4}$, and a batch size of $32$.
@@ 4.1 Datasets
2D Lemniscate Sequence. We simulate a number of $1400$ "lemniscate" series ($64^{2}\times T$), where $T+1=12$ (examples are shown in Fig. 2). For each reference image $I_{i}^{0}$, we first generate a 2D contour with its coordinates computed by $\{^{x=a\cdot\text{cos}(\alpha)/(\text{sin}^{2}(\alpha)+1)}_{y=a\cdot\text{sin}(\alpha)\text{cos}(\alpha)/(\text{sin}^{2}(\alpha)+1)}$, where $a$ is a scaling factor, and $\alpha$ is uniformly sampled from $0,2\pi]$. We then determine the contour thickness by varying the coordinates with parameters $\sigma_{x}$ and $\sigma_{y}$, i.e., $x\pm\sigma_{x},y\pm\sigma_{y}]$. The subsequent time frames $\{I_{i}^{1},\cdots,I_{i}^{T}\}$ are deformed versions of the reference image, simulated by applying affine transformations such as scaling, rotation, and translation. We split the dataset into $1000/200/200$ sequences for training, validation, and testing.
@@ 4.2 Experimental Design
[…]